OpenStudy (anonymous):

Rewrite the equation in slope-intercept form. Then identify the slope and the y-intercept of the line. x-y=6

OpenStudy (anonymous):

x-y= 6 -y =6-x divide by the -1 in y y =6+x rearrange in y =mx+b form y= x+6 slope =1 y intercept = 6

OpenStudy (anonymous):